OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

It means that it makes the equation true. 2x = 0 x = 0 is a solution. Divide both sides by 2 to get x = 0

OpenStudy (anonymous):

it means that if a number which is a solution is put in the equation the equation will get equal to 0 example y^2 + y -6 is an equation and 2 is the solution of the equation if we put 2 in the equation the answer will be zero 4+2-6=0

OpenStudy (anonymous):

A solution to an equation is a value of the variable such that when you plug that value back into the original equation, you get a true statement.

OpenStudy (anonymous):

The value of the equation does not need to be zero. You only need to have a true statement in general when plugging the value of the variable back in. For instance, the equation 2x + 5 = 3x - 2 has solution x = 7 since when you plug it in, you get the statement 19 = 19, which is true.